# Constructing equivalences between fusion categories of quantum groups and of vertex operator algebras via quantum gauge groups

### v4

@@ 1. Introduction
For all the Lie types we obtain complete identification for the tensor product structure, all the ribbon structure, and those braiding and associativity morphisms where one or two variables are the fundamental representation $V$, respectively, with Huang-Lepowsky ribbon braided tensor structure. These are parts (a), (b), (c) of Theorem 2.1.
The term weak in the setting of quasi-Hopf algebras was first introduced by Mack and Schomerus motivated by the work of Drinfeld for quasi-Hopf algebras and Drinfeld-Kohno theorem [71]. Weakness refers to a non-unital coproduct, needed to exclude the non-physical representations with zero quantum dimension, and with the need of compatibility with the fusion rules of the category arising from a suitable model in CFT, described by a quantized universal enveloping algebra of ${\mathfrak{s}l}_{2}$ at a root of unity.
Mack and Schomerus did not consider the Hopf counterpart in their theory. Weak Hopf algebras were later introduced by Böhm and Szlachanyi motivated by the work of Mack and Schomerus, but with a strictly coassociative coproduct [7]. The definition of weak Hopf algebra has become standard and understood in the sense by Böhm and Szlachanyi.
In [12], [13], and also in this paper, weak Hopf algebras are understood with a different meaning. While a different name would be helpful to avoid confusion, so far we have not been able to find a good name for our algebras to clarify the situation, and we use the term weak Hopf algebras in the sense defined in Sect. 10 in [13].
In [13] we develop a general theory of weak Hopf algebras, review and expand the theory of weak quasi-Hopf algebras, mainly Tannaka-Krein aspects, both in the algebraic and in the $C^{*}$-algebraic setting, with non-trivial unitary structure, discuss a suitable Drinfeld-Kohno theorem, and its application to the construction of semisimple weak Hopf and quasi-Hopf algebras with structure naturally compatible with that of the involved ribbon braided tensor categories, for quantum groups at roots of unity and affine vertex operator algebras at positive integer levels, including unitary structure as heredited from the unitary structure of the category due to Wenzl [99]. We follow the strategy of Drinfeld for Drinfeld-Kohno theorem which includes the comparison of braiding morphisms of the two settings, but we do not include the comparison of the associativity morphisms in our analogue, Theorem 30.4 in [13] (more details are given later). We construct a Drinfeld twist for our weak Hopf algebras useful for the braided symmetry in a way similar to Drinfeld method, by taking a square root of the coboundary matrix, in our case this is slightly more delicate because of the truncation procedure and analytic aspects. Similarly to Drinfeld-Kohno theorem, we establish equality of the associativity morphisms arising from quantum group fusion category after the twist applied to our weak Hopf algebras with Huang-Lepowsky associativity morphisms via a uniqueness theorem of a cohomological nature, but in our case this theorem seems different, in that we use cohomology of the pentagon equation on the generating representation $V$ and extensions to quantum case and Lie types of Schur-Weyl duality. We shall come back to this in the last part of the introduction and in the following sections of this paper.
We apply our concrete twist to transport unitary structures to representation categories of affine vertex operator algebras at positive integer levels with the tensor structure due to Huang and Lepowsky, for all the mentioned Lie types, by transferring the unitary structure from quantum group fusion category as constructed in [99]. On the other hand, the remaining Lie types may be either derived from Finkelberg equivalence, making now the ribbon braided fusion category $\tilde{\mathcal{O}}_{k}$ unitary, or treated with our partial results in the setting of Huang-Lepowsky structure for representation categories of vertex operator algebras. Direct constructions of unitary structures in this setting have been obtained by Gui as already mentioned [43], [44], [45], [46].
Our weak Hopf algebras while having a non-unital coproduct, are not in the class of Bohm and Szlachanyi. For example their coproduct is not strictly coassociative. They are not extension of Bohm and Szlachanyi weak Hopf algebras, and we know from results in [13] that the intersection of the two classes are the genuine Hopf algebras.
On the other hand, in our opinion the weak Hopf algebras considered in [12] and [13] still deserve to be regarded as Hopf algebras in a weak sense, for various reasons. On one hand they seem to play the role of the quantized universal enveloping algebras in Drinfeld-Kohno theorem, following the original motivation of Mack and Schomerus. Moreover, their representation category is characterized by a tensor functor to the category of Hilbert spaces with non-invertible natural transformations (projections endowed with natural inverse inclusions, in our case) satisfying tensoriality diagrams, which reduces to the notion of tensor functor when the natural transformations are invertible. Their counit is antimultiplicative and the antipode satisfy the same relations as in the case of Hopf algebras.
As already mentioned, we prove the complete identification of braiding and associativity morphisms of the two structures for all the classical Lie types $A$, $B$, $C$, $D$, and the case $G_{2}$ in part (d) of Theorem 2.1, by further refining methods of the previous part but we use more structure in Sect. 9 of [13] that has not emerged in other parts.
As above mentioned, in the type $A$ case, the extra structure that we need to use is known as quantum Schur-Weyl duality between the vector representation and the natural representation of the braid group between the tensor powers of the vector representations.
In the classical case, Schur-Weyl duality is a key property for the general linear groups starting with the work of Schur and Weyl. We note that Schur-Weyl duality also takes an important role in Doplicher-Roberts abstract duality theory for compact groups and applications to construction of the field algebra. This classical duality emergences as the starting point for the construction of global gauge symmetry groups in tensor categories arising in [24].
Duality with the braid group representation is a concept at the basis of the origin of quantum groups following V. F. R. Jones [64].
More generally, for Lie types beyond $A$, we need a corresponding duality between the centralizer algebras of the truncated tensor powers of the fundamental representation $V$ of the Drinfeld-Jimbo quantum group $U_{q}({\mathfrak{g}})$ at the root of unity and the natural representation of the braid group in these algebras, and this leads to the restriction stated in (d). This allows us to apply our abstract uniqueness Theorem 9.19 in [13] for braided tensor structures in a pre-tensor category.
In our work, the need of quantum duality between the braid group representation and the fundamental representation $V$ emerged on the occasion of the proof of our earlier Lemma 7.10, which was not complete in version 7 of [13], and we worked on completion with the same methods. It was for us a sudden realization of an even greater recognition than we had realized up to that point of the contribution of the authors on extensions of Schur-Weyl duality which has historical interest on his own in representation theory, and more work by Wenzl and his collaborators, and several other authors in the classical and quantum cases for the various Lie types. We shall come back to this in more detail.
In our setting, quantum extension of Schur-Weyl duality to the Lie types has a role to see uniqueness of the associativity morphisms in a tensor category associated to the given braided symmetry with a generating representation $V$, see Theorem 9.19 in [13]. This applies to both fusion categories associated to quantum groups at roots of unity and of affine VOAs.
Presently it is not known to the author whether the fundamental representation for the exceptional Lie types $E$ and $F$ satisfies braid group duality. The $E_{8}$ case is more delicate, because fusion of tensor powers of $U_{q}({\mathfrak{g}})$ at the root of unity is more complicated, as explained in [99], but more work by Wenzl is available for the study of the centralizer algebra [100]. We are grateful to M. Yamashita for pointing out this reference to us.
Among possible applications of our proof are the study of unitary structures of representation categories of vertex operator algebras, due to the clarity offered by a quantum group viewpoint of the same problem as explained in [99]; the study of equivalences between representation categories of conformal nets and vertex operator algebras via the construction of a common quantum gauge group; to extension of Doplicher-Roberts field algebra construction from the local observable algebra to the conformal field theory setting. Some more discussion may be found in the outlook section 13.
@@ 2. The main result
    endowed with the same structure maps as in (a) but w.r.t. the new ribbon braided tensor structure of ${\rm Rep}_{\rm QG}(A(V_{{\mathfrak{g}}_{k}}))$ described in (b), satisfies the ribbon braided tensor equivalence equations for the ribbon structure for all objects, the braid morphisms for pairs $(V_{\lambda},V)$ and $(V_{\lambda},V)$, and the associativity morphisms for triples
@@
[…]
@@
The definition of unitary coboundary weak Hopf algebra, its restricted version named unitary compatible coboundary weak Hopf algebra, are given in Sects. 31, 32 in [13], by abstracting features of the hermitian structure of $U_{q}({\mathfrak{g}})$ for $|q|=1$, discovered by Wenzl in [99] in his proof of unitarity of Kirillov inner product, for the associated fusion category as certain roots of unity, based on a complex Drinfeld coboundary matrix. The more general notion admits a categorical characterization generalizing Doplicher-Roberts symmetric functors.
For part (b), we refer the reader to Theorem 3.2 (a) in [13] and further indications therein where the statement is explained.
We would like to notice that the passages described in (2.1) are explicit. The first and the last equivalences are due respectively, to the work by Zhu combined with that of Huang and Lepowsky. The last is the functor due to Wenzl enriched with our weak tensor structure, preserving Drinfeld coboundary structure derived from the ribbon and $R$-matrix, see the above cited sections in [13].
@@
The middle equivalence in (2.1) is application of the explicit symmetry provided by our analogue of Drinfeld-Kohno theorem 30.4 in [13]. This symmetry is given by a Drinfeld twist explicitly defined by the action of the $R$-matrix, more precisely by the braided and ribbon structure in the two settings, that allow to see in the vertex operator algebra setting same structure as that in the quantum group setting.
3. Motivation and framework of our work. Constructing braided tensor equivalences via a quantum gauge group given by a discrete weak Hopf algebra and a Drinfeld twist.
Before going into the explanation of the statement and outline the main tools of the proof in more detail, we summarize the general framework that motivated and led to our work. In the following, where not indicated the cited sections and results refer to [13], and also for unexplained notation.
We were originally motivated by Doplicher-Roberts abstract duality theory for compact groups and their application to the construction of a global compact gauge group and reconstruction of the net of field algebras from the observable algebras in high dimensional quantum field theories [23], [24].
In the setting of low dimensional quantum field theories, starting with the fusion categories of quantum groups at all the roots of unity that are known to correspond to unitary structures by the work of Wenzl and Xu [99], [105], we in particular construct discrete quantum group-like structures with non-trivial twisted unitary structures for all the Lie types as in [99]. These discrete quantum groups are thought of as duals of compact quantum groups, but their preduals admit no analogy to the case of the compact quantum groups of Woronowicz, because they are non-associative algebras. However, Woronowicz work was strongly inspiring for us. Indeed our first examples in the type $A$ case were first constructed in our previous work [12] via an abstract Tannaka-Krein method in the search of a noncommutative function algebra, using the fusion rules of the corresponding fusion category.
On the other hand, their duals are quite well-behaved discrete $C^{*}$-algebras with almost associative coproducts, antimultiplicative antipode and multiplicative counit satisfying the same relation as for Hopf algebras. In this sense they resemble the duals of the compact quantum groups, the main differences being non-unitality of the coproduct and the unitary structure of the tensor product of their representations, which is twisted via a positive operator associated to the $R$-matrix (Drinfeld coboundary matrix), and makes the representation of the braid group unitary. The need of a twist for the tensor product representation arises from the fact that coproduct and -involution satisfy an anticommutativity condition, rather than commutativity as it usually is the case for Hopf algebras in the operator algebra setting. This anticommutativity is derived from the same property of the -involution of the non-semisimple $U_{q}({\mathfrak{g}})$, first pointed out by Wenzl [99].
The weak Hopf algebras constructed in [12] have been extended to all Lie types in Sections 31, 32 of [13] with uniform methods for all the Lie types.
The construction of associativity morphisms satisfying the two hexagonal equations with respect to natural braiding morphisms associated to representations of the universal enveloping algebra $U({\mathfrak{g}})[[h]]$ of a simple Lie algebra (in a formal setting) with classical coproduct, were first solved by Drinfeld with the introduction and construction of quasi-Hopf algebras, via the Knizhnik-Zamolodchikov differential equation. This category is known as Drinfeld category, motivated by previous work by Kohno on monodromy of a connection associated to the vector representation for all ${\mathfrak{g}}$ of classical type [69]. Drinfeld developed the remarkable and beautiful Drinfeld-Kohno theorem connecting $U({\mathfrak{g}})[[h]]$ with its natural $R$-matrix and non-trivial co-associator with the quantized universal enveloping algebra $U_{h}({\mathfrak{g}})$ via a Drinfeld twist, describing a ribbon braided tensor equivalence. In this case, rigidity of Drinfeld category is derived from the equivalence constructed by Drinfeld in Drinfeld-Kohno theorem [25], [27], [28].
Our algebras are examples of Mack-Schomerus weak quasi-Hopf algebras introduced in [71] as generalization of Drinfeld quasi-Hopf algebras with non-unital coproduct. For ${\mathfrak{g}}={\mathfrak{s}l}_{2}$ our examples reduce to Mack-Schomerus examples for the algebraic structure. We understand their lack of coassociativity as quite mild, because their representation category is endowed with a weak unitary tensor functor in the category of Hilbert spaces. Thus we regard them as the Hopf counterpart of weak quasi-Hopf algebras, in a position similar to that of Hopf algebras among quasi-Hopf algebras, that in our knowledge has been overlooked.
Here the term weak is understood in the sense first introduced by Mack and Schomerus, the coproduct is non-unital, a feature needed to exclude the non-physical representations with zero quantum dimension, and with the need of compatibility with the tensor structure of the category. The weakness of tensoriality of the embedding functor of their representation category to the category of Hilbert spaces means that the structure natural transformations are surjections onto subspaces of the tensor product spaces endowed with left inverses, rather than isomorphisms, and satisfy two tensoriality diagrams.
Much of Woronowicz theory can be extended to our case. From this via a quantization isomorphism described by Wenzl in [99] and the construction of a Drinfeld twist we construct and describe in some more detail later a quantum group-like structure on the Zhu algebra associated to the affine models described as in the statement of the main result.
These constructions are in some quantum analogy with Doplicher-Roberts construction in the setting of high dimensional algebraic quantum field theory, in that the categories that we study are not given as Tannakian categories. In our case, differently from their case, we have naturally defined linear functors emerging as the minimum energy functor due to Zhu in the setting of vertex operator algebras. Also differently form their case, where the linear functor needed to be constructed but the tensor structure was trivial, in our case we need to construct tensor structures associated to these naturally occurring linear functors to the category of vector spaces.
Our starting point is what we shall refer to as Wenzl functor associated to the quantum group fusion category described in [99]. This paper plays a central role, both on the side of quantum groups and on the side of affine vertex operator algebras, as we next try to explain.
As said, on the the side of quantum groups, our approach has been influenced by results in quantum group theory by Woronowicz but also by the dual viewpoint by Majid for quasi-Hopf algebras, extended to weak quasi-Hopf algebras by Haring-Oldenburg. We start reviewing these Tannakian reconstruction theorems from the viewpoint of Majid and Haring-Oldenburg in Sect. 6, 7, in [13], where the reader may find indication of the original references. Moreover Kazhdan and Wenzl theory is also important to us, and the reader may consult Sect. 25 in [13] and references therein. It is by no means an exaggeration to assert that our paper [13] would have not existed without the paper by Neshveyev, Tuset on Kazhdan-Lusztig theorem for compact quantum groups [78]. This paper was strongly inspiring for our strategy, although methods are different. For similar reasons, we were also guided by the presentation of Drinfeld-Kohno theorem in the book by Bakalov and Kirillov [1]. These last two references first inspired our work in the search of a unified treatment following a Tannakian approach in the idea of Majid, in our case. An important aspect in our case as far as Majid Tannakian duality for quasi-Hopf algebras, is the search of natural tensor structures for the involved embedding functors into the category of Hilbert spaces (Wenzl functor and Zhu functor in the two settings).
Primary fields first emerged in the physics literature, [5], using previous work by Feigin and Fuks [32] and Kac [65] on the representation theory of the Virasoro algebra. In 1984 Witten introduced what are now known the WZW models, due to Wess, Zumino, Novikov and Witten [103]. In the same year, Knizhnik and Zamolodchikov showed that their correlation functions satisfy a system of partial differential equations [68].
Tsuchiya and Kanie started a rigorous study in [95]. Goddard, Kent and Olive realized the chiral unitary Virasoro minimal models via a coset construction [40].
Moore and Seiberg conjectured some analytic properties of the quantum fields and described the construction of a modular category associated to a general conformal field theory.
In the same paper [69] Kohno had also considered monodromy of $n$-point functions of a conformal field theory on ${\mathbb{P}}^{1}$ associated to the affine Lie algebra $\hat{\mathfrak{s}l}_{2}$.
The mathematical contribution by Tsuchiya and Kanie [95], and the subsequent works by A. Wassermann [98] and Toledano-Laredo [94] on affine Lie algebras, compared with the structure that had previously emphasized by Wenzl in the study ofquantum group fusion categories and their unitary structures [99] that led to our approach. The reader may find reference to more of the several original papers in mathematics and physics in Sections 34–39 in [13].
4. The general plan of our proof.
We briefly explain here the plan of our proof of the main theorem for parts (a), (b), (c), and (d). In the following sections we describe the main ideas to prove these parts in greater detail.
Part (a) is an application of a general construction of a natural structure of a (ribbon, braided) tensor equivalence on a given equivalence ${\mathcal{C}}\to{\rm Rep}(A)$ of semisimple linear categories endowed with an inverse linear equivalence ${\rm Rep}(A)\to{\mathcal{C}}$, with $A$ a discrete algebra. We outline the main ideas of this tensor structure construction in the next section. The reader may refer to Section 16 in [13] for the detailed construction.

### v5

@@ 1.6. Main theorem
    For the remaining exceptional Lie types we obtain the identification for the tensor product structure, the ribbon structure, and those braiding and associativity axioms for morphisms where one or two variables are the fundamental representation $V$, respectively.
The previous result is stated in greater detail as Theorem 3.1. Our starting point for this theorem was in the context of constructive rigorous conformal field theory, in the setting of operator algebras. The main references important for us to start to set up the connection between quantum groups and affine vertex operator algebras, are the work of Wenzl [99], and the work by Wassermann [98] in the setting of affine Lie algebras. A key concept is played by the notion of primary field, and its initial term, which determines the primary field uniquely, as shown by Tsuchiya and Kanie [95]. We use this fact in a fundamental way to connect quantum group fusion category to representations of the Zhu algebra and their special tensor products with Wenzl fundamental representation, see Sect. 35 in [10]. Most importantly, we describe a connection between Wenzl and Wassermann work via Wenzl quantisation curve, and building on their work in [10] we obtain Theorems 2.1, 2.2, Cor. 2.3, and Theorem 2.4 which compares with the tensor product structure by Huang and Lepowsky, to which we refer the reader also for important historical references. The connection from affine Lie algebra setting to Huang-Lepowsky tensor product theory may be found in Sections 33–38 in [10].
A main idea that we follow is to consider the fundamental representation $V$ of $\mathfrak{g}$} for each Lie type as in [99] where the author explains its generating property and good fusion structure with respect to tensor products with an arbitrary irreducible representation allowing a well-defined truncation procedure to obtain non-negligible objects, with the aim of studying unitarizability of the fusion category.
As anticipated in the abstract, the restriction to the mentioned Lie types is due to the fact that we need to use a key property of the theory of quantum groups, namely the extension to the quantum case of generalized Schur-Weyl duality. Currently, this property is known to hold for these Lie types. This is part (d) of the same theorem.
The final stage of the proof, detailed in Sect. 9–11, relies on a process of cohomological reduction and heavily utilizes extensions of quantum Schur-Weyl duality. By explicitly exploiting the generating property of the braid group representation within the centralizer algebras of the truncated tensor powers of the fundamental representation, we achieve the direct equivalence with Huang-Lepowsky braided tensor structure of ${\rm Rep}(V_{{\mathfrak{g}}_{k}})$ necessary to solve Huang’s problem for the classical and $G_{2}$ Lie types.
The paper is organized as follows. In Sect. 2 we give the main definition of unitary coboundary weak Hopf algebra and recall the construction of $A_{W}({\mathfrak{g}},q)$. In Sect. 3 we state the main result, Theorem 3.1. Sects 4–7 explain the main ideas to prove parts (a)–(d) of this result. We give details for the completion of the proof of part (d) in Sects 9–11. Sect. 12 concerns connection with Doplicher-Roberts theorem regarding the role of generalized Schur-Weyl duality. Sect. 13 concerns future developments. We put in the Appendix historical remarks on the proof of rigidity in module categories of affine Lie algebras and affine vertex operator algebras at positive integer level.
2. Unitary coboundary weak Hopf algebras $A_{W}({\mathfrak{g}},q)$
The notion of weak quasi-Hopf algebra was first introduced by Mack and Schomerus motivated by the work of Drinfeld on quasi-Hopf algebras, specifically the Drinfeld-Kohno theorem [70].
The weakness refers to a non-unital coproduct, and is needed to construct a coalgebra structure on the quotient $A$ of $U_{q}({\mathfrak{g}})$ with $q$ a root of unity by the ideal of elements annihilated by representations with zero quantum dimension. Since the ideal is not a coideal, the quotient algebra $A$ is not a bialgebra in the natural way. For ${\mathfrak{s}l}_{2}$, Mack and Schomerus defined a weak coproduct on $A$ and an associator derived from the fusion rules of the semisimple fusion category ${\mathcal{C}}({\mathfrak{s}l}_{2},q)$, a counit and a natural antipode derived from $U_{q}({\mathfrak{s}l}_{2})$, and have verified the axioms of a weak quasi-Hopf algebra on their example. Their construction is based on the special multiplicity free fusion rules of the category. In [10] we have extended Mack and Schomerus construction to all Lie types and to all roots of unity of the unitary type–i.e. the roots of unity appearing in the statement of our main Theorem 1.6–and we have derived algebraic and analytic properties that were overlooked in the literature.
The main algebraic property that our quotient algebra $A_{W}({\mathfrak{g}},q)$ of $U_{q}({\mathfrak{g}})$ satisfies is that of being the Hopf counterpart of a weak quasi-Hopf algebra, in a position completely analogous to Hopf algebras among quasi-Hopf algebras. For this reason, we shall refer to $A_{W}({\mathfrak{g}},q)$ as a weak Hopf algebra. We have established the weak Hopf algebra structure of $A_{W}({\mathfrak{g}},q)$ building upon the quantum group fusion category ${\mathcal{C}}({\mathfrak{g}},q)$ and Wenzl’s linear fibre functor $W$ associated to it [99].
Here we limit ourselves to recall the axioms of the definition of weak Hopf algebra which mostly differ from those of Hopf algebras. For the complete properties of antipode, counit and $R$-matrix of weak Hopf algebras we refer the reader to Sects. 9–14 of [10].
Definition 2.1.
Let $A$ be a semisimple unital complex algebra endowed with a coproduct $\Delta:A\to A\otimes A$ given by a not necessarily unital homomorphism and a counit satisfying the usual counit axioms for bialgebras. The tensor product $\rho\otimes\sigma$ of two f.d. representations of $A$ is the representation defined by $\Delta$ restricted to the range of $\Delta(I)$ in the tensor product space.
We call $A$ a weak bialgebra if the category ${\rm Rep}(A)$ of representations of $A$ becomes a tensor category with associativity morphisms
$\alpha_{\rho,\sigma,\tau}:(\rho\otimes\sigma)\otimes\tau\to\rho\otimes(\sigma\otimes\tau),\quad\rho,\sigma,\tau\in{\rm Rep}(A)$
defined by the action of $1\otimes\Delta(\Delta(I))\Delta\otimes 1(\Delta(I))$ between the representation spaces, with $I$ the unit of the algebra, and $1$ the identity map. In particular, if the coproduct is unital, $A$ is a coassociative bialgebra.
A weak Hopf algebra is a weak bialgebra endowed with an antipode $S:A\to A$ satisfying the axioms of a Hopf algebra antipode.
Among weak quasi-bialgebras, weak bialgebras admit a characterization in terms of weak tensoriality of their fibre functor.
Proposition 2.2.
A semisimple weak quasi-bialgebra $A$ is a weak bialgebra if and only if the fibre functor ${\mathcal{F}}:{\rm Rep}(A)\to{\rm Vec}$ is a weak tensor functor. By this we mean that the associated natural transformations given by projection and inclusion, respectively:
$F_{\rho,\sigma}:{\mathcal{F}}(\rho)\otimes{\mathcal{F}}(\sigma)\to{\mathcal{F}}({\rho\otimes\sigma})$
and
$G_{\rho,\sigma}:{\mathcal{F}}({\rho\otimes\sigma})\to{\mathcal{F}}(\rho)\otimes{\mathcal{F}}(\sigma)$
satisfy for all objects $\rho$, $\sigma$, $\tau$,
(2.1) $F_{\rho,\sigma}\circ G_{\rho,\sigma}=1_{{\mathcal{F}}({\rho\otimes\sigma})}$
(2.2) $F_{\rho^{\prime},\sigma^{\prime}}\circ{\mathcal{F}}(S)\otimes{\mathcal{F}}(T)={\mathcal{F}}(S\otimes T)\circ F_{\rho,\sigma},\quad{\mathcal{F}}(S)\otimes{\mathcal{F}}(T)\circ G_{\rho,\sigma}=G_{\rho^{\prime},\sigma^{\prime}}\circ{\mathcal{F}}(S\otimes T)$
for objects $\rho$, $\sigma$, $\rho^{\prime}$, $\sigma^{\prime}\in{\mathcal{C}}$ and morphisms $S:\rho\to\rho^{\prime}$, $T:\sigma\to\sigma^{\prime}$.
(2.3) ${\mathcal{F}}(\alpha_{\rho,\sigma,\tau})=F_{\rho,\sigma\otimes\tau}\circ 1_{{\mathcal{F}}(\rho)}\otimes F_{\sigma,\tau}\circ\alpha^{\prime}_{{\mathcal{F}}(\rho),{\mathcal{F}}(\sigma),{\mathcal{F}}(\tau)}\circ G_{\rho,\sigma}\otimes 1_{{\mathcal{F}}(\tau)}\circ G_{\rho\otimes\sigma,\tau}$
(2.4) ${\mathcal{F}}(\alpha_{\rho,\sigma,\tau}^{-1})=F_{\rho\otimes\sigma,\tau}\circ F_{\rho,\sigma}\otimes 1_{{\mathcal{F}}(\tau)}\circ{\alpha^{\prime}}^{-1}_{{\mathcal{F}}(\rho),{\mathcal{F}}(\sigma),{\mathcal{F}}(\tau)}\circ 1_{{\mathcal{F}}(\rho)}\otimes G_{\sigma,\tau}\circ G_{\rho,\sigma\otimes\tau},$
where $\alpha$ are the associativity morphisms of ${\rm Rep}(A)$ defined by the associator of $A$, and $\alpha^{\prime}$ are the (trivial) associativity morphisms of $\rm Vec$}.
We refer to Sect. 6, 7, 10 in [10] for the general definition of weak quasi-Hopf algebra, their $R$-matrix how to obtain a rigid braided tensor category from it. This structure of weak quasi-Hopf algebra is stable under a twist operation–first introduced by Drinfeld for quasi-Hopf algebras–. The twist of a weak quasi-Hopf algebra is automatically a weak quasi-Hopf algebra, and provides an equivalent rigid tensor category.
We next describe the properties concerning the unitary structure of weak quasi-Hopf algebras that are most relevant for this paper.
Definition 2.3.
A (compatible) unitary coboundary weak quasi-Hopf algebra $A$ is defined by the following data:
  • a)
    A unital weak quasi-Hopf algebra $A$ endowed with a $C^{*}$-algebra involution with an antipode $(S,\alpha,\beta)$
  • b)
    a ribbon structure $(R,v)$ for $A$ associated to $(S,\alpha,\beta)$ such that the ribbon element $v\in A$ is unitary,
  • c)
    a unitary central square root $w\in A$ of $v$ such that $\varepsilon(w)=1$, $S(w)=w$,
  • d)
    $\tilde{A}=A^{{\rm op}}$ as quasitriangular weak quasi-bialgebras,
  • e)
    $\overline{R}=Rw^{-1}\otimes w^{-1}\Delta(w)$ is positive.
The element $R$ is the $R$-matrix; $v\in A$ is a central element defining the ribbon structure; the operator $\overline{R}$ is Drinfeld coboundary matrix and gives the $\Omega$-involution $\Omega=\overline{R}$ in the sense of Sect. 11, 12 in [10]. The $\Omega$-involution induces a -involution on ${\rm Rep}(A)$ making it into a tensor -category. In the case where the -involution is induced by a coboundary matrix, the braided symmetry of ${\rm Rep}(A)$ induced by the $\overline{R}$-matrix is unitary by Theorem 27.9 in [10].
The triple $(S,\alpha,\beta)$ describes the antipode of a weak quasi-Hopf algebra, and makes ${\rm Rep}(A)$ into a rigid category. The presence of invertible elements $\alpha$, $\beta$ is necessary in the quasi-cases, it arises from twist deformations of the usual antipode of a Hopf algebra. If $A$ is weak Hopf algebra, its antipode is required to satisfy the same axioms as a Hopf algebra antipode: is defined with $\alpha=\beta=I$; $\varepsilon$ denotes the counit and satisfies the same axioms as in the Hopf algebra case; $A^{{\rm op}}$ and $\tilde{A}$, denote the opposite and adjoint weak quasi-Hopf algebras. The opposite structure $A^{{\rm op}}$ is widely used in the literature, while the adjoint structure $\tilde{A}$ is defined by coproduct $\tilde{\Delta}(a)=\Delta(a^{*})^{*}$, and similarly for the remaining structure, see Sect. 10 in [10].
Except for the $C^{*}$-structure of the algebra and the positivity requirement of the coboundary operator in e), the axioms are satisfied by the -algebra of $U_{q}({\mathfrak{g}})$, which for this reason represents the motivating example of our definition. These properties of $U_{q}({\mathfrak{g}})$ and positivity of a truncated coboundary operator operator of $U_{q}({\mathfrak{g}})$ on certain special tensor products are due to Wenzl [99], see also Sect. 30 in [10].
Compatible unitary weak quasi-Hopf algebras in the sense of Definition 2.3 have been introduced and studied in Sect. 29 of [10]. They satisfy the stronger property d) than their weaker variant considered in Sect. 27 of the same paper, where d) is relaxed up to a so called trivial twist. If the coproduct is unital, the two definitions coincide. We do not have examples in the weak case that distinguishes among the two definitions.
The weaker variant is interesting for the fact that it admits a categorical characterization extending the notion of symmetric fibre functor –a core notion in Doplicher-Roberts theory.– to the coboundary symmetry induced by $\overline{R}$ in ${\rm Rep}(A)$. The stronger notion on the other hand allows a formulation of an abstract Drinfeld-Kohno theorem, see Theorem 29.4 in [10], that we use in our main theorem. In this paper we shall not need the weaker version, therefore we shall always assume the compatibility condition given by property d) when not clearly stated.
In [99] Wenzl has constructed a linear -functor
$W:{\mathcal{C}}({\mathfrak{g}},q)\to{\rm Hilb}.$
The following result has been shown in Theorem 31.16 in [10].
Theorem 2.4.
The semisimple -algebra $A_{W}({\mathfrak{g}},q):={\rm Nat}(W)$ of natural transformations of $W$ admits a natural structure of unitary compatible coboundary weak Hopf algebra derived from ${\mathcal{C}}({\mathfrak{g}},q)$ and $W$.
Thanks to Theorem 29.4 in [10]–our version of Drinfeld-Kohno theorem–the Zhu algebras $A(V_{{\mathfrak{g}}_{k}})$ become unitary coboundary weak quasi-Hopf algebras with a Drinfeld twist from $A_{W}({\mathfrak{g}},q)$, by our version of Drinfeld-Kohno theorem. This is summarized in the following section, and detailed in Sect. 33 of [10].
@@ 3. The main result
    endowed with the same natural transformation as in (a) but w.r.t. the new ribbon braided tensor structure of ${\rm Rep}_{\rm QG}(A(V_{{\mathfrak{g}}_{k}}))$ described in (b), satisfies the ribbon braided tensor equivalence equations for the ribbon structure for all objects, the braid morphisms for pairs $(V_{\lambda},V)$ and $(V_{\lambda},V)$, and the associativity morphisms for triples
@@
    (3.1) ${\rm Rep}_{\rm HL}(V_{{\mathfrak{g}}_{k}})\xrightarrow{Z}{\rm Rep}_{\rm QG}(A(V_{{\mathfrak{g}}_{k}}))\xrightarrow{}{\rm Rep}(A_{W}({\mathfrak{g}},q))\xrightarrow{{\rm TK\ ribbon\ equiv}}{\mathcal{C}}({\mathfrak{g}},q)$
@@
We would like to notice that the equivalences described in (3.1) are explicit. The equivalence on the left is Zhu’s linear equivalence endowed with the natural transformations given by the restriction to the zero modes the braided tensor structure by Huang and Lepowsky. The fact that the braided tensor structure of the range of $Z$ induced by Huang-Lepowsky structure in (a) coincides with the structure arising from quantum groups on the Zhu algebra, follows from our work in (c), done in [10], and the first statement of part (d), established in this paper. The equivalence on the right is the functor due to Wenzl enriched with our weak tensor structure by Theorem 2.4. The middle equivalence is an application of the explicit symmetry provided by our analogue of Drinfeld-Kohno theorem 29.4 in [10], and stated in parts (a) and (b) of Theorem 2.2 in [10]. This symmetry is given by a Drinfeld twist explicitly defined by the action of the $R$-matrix, more precisely by the braided and ribbon structure in the two settings, that allow to see in the vertex operator algebra setting same structure as that in the quantum group setting.
4. The general plan of our proof
Part (a) of Theorem 3.1 follows from the application to Zhu’s functor of the general construction. Let us consider a (ribbon, braided) semsimple tensor category $\mathcal{C}$} together with a discrete complex algebra $A$ (a direct sum of full matrix algebras), a linear equivalence ${\mathcal{C}}\to{\rm Rep}(A)$ and an inverse linear equivalence ${\rm Rep}(A)\to{\mathcal{C}}$. We wish to use these equivalences to transfer all the structure of $\mathcal{C}$} to ${\rm Rep}(A)$, and make it into a (ribbon, braided) tensor category in such a way that the given equivalences become (ribbon, braided) tensor equivalences. The definition of the structure on ${\rm Rep}(A)$ is very natural and verification of the axioms of (ribbon, braided) tensor category and the natural transformations on the equivalences and the needed properties is straightforward. The special case of making ${\rm Rep}(A(V_{{\mathfrak{g}}_{k}}))$ into a ribbon braided tensor category from Huang-Lepowsky structure on ${\rm Rep}(V_{{\mathfrak{g}}_{k}})$ corresponds to pass from the structure on a category of infinite dimensional modules to the category of their zero modes. We have explicitly described this construction in Sect. 37 in [10]. This construction was also considered in [72]. We have used a similar construction, in the reverse direction, and for a different purpose in Theorems 15.6 and 22.10 of [10].